OpenStudy (anonymous):

not sure if i got this right. but i got 1 \[\log_{6} (x+3)+\log_{6} (x-2)=1\]

mathslover (mathslover):

\[\log_6(x+3)+log_6(x-2)=log_6(x+3)(x-2)\] were there any values given for x ?

OpenStudy (anonymous):

the only equation i was given is this:\[\log_{6} (x+3)+\log_{6} (x-2)=1\]

mathslover (mathslover):

ok so u have to prove this ?

OpenStudy (anonymous):

i just have to solve it.

OpenStudy (anonymous):

use log properties

OpenStudy (anonymous):

to combine into simpler equation

OpenStudy (anonymous):

the answer i ended up with was 1 but I'm not sure if i got it right

mathslover (mathslover):

yes ! : \[\log_a(c)+log_a(b)=log_a(bc)\]

OpenStudy (anonymous):

did you get x^2+x-7=0 as your quadratic?

mathslover (mathslover):

how will u get 0 right up there ?

mathslover (mathslover):

http://www.wolframalpha.com/input/?i=log_6(x%2B3)%2Blog_6(x-2)&t=crmtb01 there does not seem to value coded as 1

OpenStudy (anonymous):

lol its log base 6

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

so what i did was combine them and got log (base 6) x^2+x-6)=1

OpenStudy (anonymous):

x^2+x-12=0?

OpenStudy (anonymous):

and solve x

OpenStudy (anonymous):

and then plug x back into original equation

OpenStudy (anonymous):

nvm! i mixed the signs up and just realized what i did! i didn't get 12 and i got 0

OpenStudy (anonymous):

excluding values that makes log negative

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

so if you got 1 then ur answer is wrong

OpenStudy (anonymous):

yeah i know. i realized what i did wrong.

OpenStudy (anonymous):

log(1-2)=log(-1) which is not true

OpenStudy (anonymous):

i got 4 for my answer